# PLM-Based Discrete Diffusion Language Models with Entropy-Adaptive Gibbs Sampling

### v2

@@ 3.3 Training and Inference
             $x^{t-1}\leftarrow\text{Sampling}(x^{t},l^{*},Y)$ // Sample from the previous timestep
       end if
end for
Algorithm 1 ENS and EAGS Algorithm
Our whole process is in Algorithm 1. Distinct from the prevailing methodologies in diffusion models as described by Ho et al. (2020a) and Austin et al. (2023), we do not employ the PLM parameterization strategy $\widetilde{p}_{\theta}(\widetilde{z}_{0}|z_{t},t)$, which preserves the original semantic embedding spaces during the training phase. We empirically find that PLM parameterization restricts the diversity of generated responses. Instead, to ensure the completeness of sentences, we implement Minimum Bayes Risk (MBR) decoding in the final generation step. We follow the traditional diffusion loss in Equation (Ho et al., 2020b), changing MSE with Cross Entrpy Loss.
4 Experiments
4.1 Tasks and Datasets
Social Datasets The Paradetox dataset (Logacheva et al., 2022) focuses on removing profanities. The Deontology from ETHICS dataset (Hendrycks et al., 2023) evaluates ethical judgment based on scenarios. Paraphrase The QQP dataset (Wang et al., 2017) assesses paraphrase detection by treating one question as a paraphrase of another. Question Generation We use the Quasar-T dataset (Dhingra et al., 2017) to generate questions from passages and answers. Open-ended Generation The RocStories (Mostafazadeh et al., 2016) dataset is used to generate coherent story continuations from a given context. Detailed explanations are available in Appendix C and Dataset statistics are shown in Table 8.
4.2 Baselines
Since our objective is centered on generation tasks, we compare Diffusion-EAGS with four categories of baselines with the model similar to the size of RoBERTa-base (Liu et al., 2020): Auto-regressive Models (ARMs), CMLMs (Conditional Masked Language Models, Ghazvininejad et al. (2019a)), CDLMs, and DDLMs. For ARMs (Vaswani et al., 2023), we employ GPT-2 (Radford et al., 2019), which is renowned for its proficiency in generation tasks. For CMLMs, we utilize CMLM (Ghazvininejad et al., 2019a) and DisCo (Kasai et al., 2020), which are transformer-based NAR models. Inference is carried out using the Mask-Predict (Ghazvininejad et al., 2019a) and the Easy-First (Kasai et al., 2020) inference algorithm respectively. For CDLMs, our baseline includes DiffuSeq (Gong et al., 2023a) and DINOISER (Ye et al., 2024). DiffuSeq is a diffusion model specifically designed for sequence-to-sequence applications, whereas DINOISER adaptively determines the range of sampled noise scales during training. For DDLMs, we utilize DiffusionBERT (He et al., 2022) , a state-of-the-art model in the research series of DiffusionLM (Li et al., 2022), and SEDD (Lou et al., 2024), a state-of-the-art model of open-ended generation in diffusion language generation domain. For SEDD, we download the pretrained version and fine-tune on it. Experimental results on LLMs can be found in Appendix E.
4.3 Metrics
Quality metrics To measure the quality of the generated texts, we use Perplexity based-on GPT-2 Large and GPT-2 XL, SOME (Yoshimura et al., 2020), grammar metric based on the neural net, LLM-c (Lin and Chen, 2023) to measure the plausibility of the narratives, LLM-t (Koh et al., 2024) to measure toxicity, and MAUVE Pillutla et al. (2021), measuring the gap between text distributions despite divergence.
Diversity Metrics Traditional diversity metrics Self-BLEU Zhu et al. (2018) and distinct-n Li et al. (2015) are employed to evaluate the generated texts. We also adopt Vendi Score (VS) Friedman and Dieng (2023), an interpretable diversity metric, which quantifies the effective number of unique samples in a given set. Both the n-gram and embedding variations are utilized, where embedding VS interprets diversity of semantics.
4.4 Experimental Details
@@
We employ roberta-base as encoder-based PLM. The learning rate is set at 5e-4, and a naive categorical sampling strategy is adopted. Adapted to data statistics, the maximum lengths for QG, QQP, and Paradetox is set to 64, while for Deontology set to 48. Furthermore, in accordance with the minimum construction length, the number of steps is configured to 5 and 20 size of MBR. We use 1 A100 GPU, and the batch size is set to 256.
@@
5 Results
As shown in Table 234, our model consistently exhibits exceptional performance in terms of text quality while simultaneously maintaining diversity when compared to baseline models.
In Table 2 Paradetox, our model demonstrates superior performance across all evaluated metrics. Such phenomenon represents that our model based on encoder-based PLMs show robustness on diverse perturbations from daily dialogues. When PPL exceeds 600, the model is considered to have failed in generating natural sequences, and is thus represented in gray color. In the context of Deontology, our model exceeds the baseline PPL and MAUVE scores, whereas SOME score represent the sufficient quality of text with the highest diversity score of 4.755. In the paradetox experiment, the text quality produced by the CMLM and SEDD models was found to be low. Consequently, these models were excluded from subsequent experiments. As illustrated in Table 3, Diffusion-EAGS generates the responses with the highest PPL score for QG, and highest MAUVE and PPL score for QQP. Additionally, in Table 4 where semantic consistency is crucial, our model significantly outperforms the original dataset’s PPL by 23 points by leveraging PLMs while maintaining a high diversity score of 4.837.
Diffusion-EAGS demonstrates the highest MAUVE score of 0.733 in Table 2-ParaDetox, and high level of text quality surpassing that of GPT-2 in Table 3 in text quality. ParaDetox is colloquial dataset including slang, numerous abbreviations, and various perturbations, so our model demonstrate robustness to such perturbations with 69.5 PPL. As for diversity, our model consistently outperforms GPT models in VS(ngram) and VS(emb) in Table 23, and 4. These results underscore our model’s efficacy in generating diverse responses.
Model PPL $\downarrow$ SOME $\uparrow$ LLM-c $\uparrow$ VS(ngram) $\uparrow$ self-bleu $\downarrow$
Original Data 100.6 0.895 1
GPT-2 88.5 0.856 0.88 4.722 0.124
DiffusionBert 318.2 0.783 0.72 4.735 0.088
SEDD 273.2 0.827 0.59 4.859 0.044
Diffusion-EAGS 67.3 0.844 0.87 4.837 0.058
Table 4: ROC Generation – Because ROC is open-ended generation task, we include SEDD in this experiment.
@@ 5 Results
DiffusionBERT exhibits limitations in text quality, as shown in Table 234. On the other hand, our model achieves significantly higher scores across all quality metrics. This observation suggests that the naive application of BERT (Devlin et al., 2019) within the diffusion process fails to fully harness the capabilities of PLMs. SEDD, an open-ended generation model shows low performance under text generation in highly constrained tasks such as ParaDetox. Additionally, our model does not require more than 5 diffusion steps, as it can adaptively recover the absorbing state through an entropy-based generation approach. A detailed comparison is provided in Appendix D.
6 Analysis
6.1 Ablation Study
To explore the effectiveness of our model’s components, we conduct ablation studies focusing on three key elements: Entropy Adaptation, stepwise Gibbs Sampling, and PLM in Table 5.
The omission of EAGS initially leads to a substantial decline in performance and text quality, producing degenerated results similar to those of traditional CMLMs.
Dataset PPL MAUVE SOME VS(ngram) VS(emb)
Diffusion-EAGS Deont 55.1 0.412 0.835 4.898 4.009
Roc 67.3 0.87 0.844 4.837 3.999
w/o EAGS Deont 667.9 0.022 0.617 4.767 3.928
Roc 1084.9 0.035 0.613 4.874 3.957
w/o Gibbs Sampling Deont 1426.7 0.011 0.584 2.378 1.923
Roc 1293.1 0.010 0.534 1.531 1.338
w/o PLM Deont $\geq$2K 0.005 0.645 4.758 3.402
Roc $\geq$2K 0.004 0.604 4.315 2.994
Table 5: Ablation Study : Deont denotes Deontology
We find that EAGS significantly contributes to a gradual entropy decrease, as shown in Appendix J. These observations highlight the critical role of selecting the sampling position based on given information in sequence generation. Additionally, excluding the use of the diffusion generation process, without the stepwise sampling, leads to a drastic reduction in overall performance, with PPL increasing by more than 1000 points and diversity scores dropping below 2. Consequently, our cMRF approach, integrated with the diffusion model, proves indispensable in aiding the PLM in effectively generating sequences. Furthermore, it is evident that without the incorporation of PLMs, the generation of natural sequences is unachievable.
6.2 Case Study: Keyword Based Generation
Our model operating within discrete space enables us to manipulate the output sequences using explicit instructions. To further explore this capability, we conduct the generation of sequences based on keywords positioned in the middle and at the end of masked sequences, which is challenging for AR models (Keskar et al., 2019). Initially, we provide the same contextual input while varying the keywords. In the masked states, we randomly select positions, replacing them with the specified keywords. The results in Table 7 demonstrate that the generated sequences seamlessly integrate the keywords with context-specific semantics.
These findings show that our model can effectively integrate specific conditional information in an interpretable manner. Experimental results also indicates its substantial potential and suitability for diverse applications where direct controllability is crucial, such as in story generation.
6.3 Bilinguality & Low Resource Settings
@@
He cried a lot over it. His parents decided to replace it for him.
Keyword not stop Jake just could not stop crying.
Jake feel It made Jake feel So much better.
would enjoy Jake said he would enjoy the new toy
… He thought the food and culture in Sofia were the best.
Keyword Bulgaria! Things were looking great in Bulgaria!
She decided to book a week long trip. She left the next day for her trip.
Keyword her trip She spent almost a week there during her trip.
Table 7: Keyword-based Generation
[…]
Although our baseline diffusion models demonstrated significant performance improvements over CMLMs, CMLMs were primarily developed to enhance translation performance. Therefore, we conducted additional experiments to further validate our methodology such as Mask-and-Predict and Easy-First, diffusion models such as Diffuseq-v2 (Gong et al., 2023b) and SEDD, traditional translation models such as mBART-50 (Tang et al., 2020) and NLLB. For evaluation metrics, we utilize sacreBLEU (Post, 2018) and neural-net scores such as COMET (Rei et al., 2020) and XCOMET (Guerreiro et al., 2023).
Table 6 shows that predicting the target sequence without leveraging a multilingual PLM proves to be challenging. All diffusion baseline models struggle to produce correct outputs. For example, the pretrained SEDD model fails to effectively leverage conditional information, even after finetuning on the training datasets, consistent with the limitation observed in Section D.3. Similar challenges arise in NAR transformer baselines. Despite constructing the vocabulary using the pretrained mBART-50 model (DisCo-m), the underlying issues remain. On the other hand, our proposed model, by incorporating a PLM, demonstrates promising results. In addition, these results show that our methodology can also be used with encoder-decoder models.
Interestingly, the output of the pretrained NLLB model (NLLB-naive-600M, not finetuned) reveal that neural network-based metrics are susceptible to the interference problem, specifically translated by other languages, even though we provide the language specific token. While such issues result in lower BLEU scores, COMET and XCOMET often interpret them as semantically coherent, indicating a potential direction for future work to improve translation evaluation metrics. Despite these phenomena, a performance gap between translation models and DDLM remains. This suggests that future research should address the semantic capabilities of diffusion models to help bridge this gap. Details are in Appendix I.
6.4 Diversity Saturation
@@
Inspired by the observation that Diffusion-EAGS consistently excel in terms of diversity across all results, we delve further into the diversity capabilities of our model. We assess the diversity performance in conditional generation compared to LLMs. We measure the VS for 5 to 100 generations under a single condition. Such experiment demonstrates the extent to which the model’s output diversity saturates, enabling a comparison of asymptotic diversity performance. The experiment is conducted on the ‘deontology’ dataset which allows high output diversity in its settings. Details of using LLMs are provided in Appendix G.
@@
7 Related Works
Efforts to integrate generative flow models into sequence generation exploit the distribution shift from a source language to a target language through a series of invertible linear transformations (Ma et al., 2019; Zhang et al., 2024). However, as DDPM (Ho et al., 2020a) demonstrate the effectiveness of generating images, diffusion models have been a major topic of interest within the field of generative flow models (Song et al., 2021a, b). To apply such diffusion methodologies to NLP, there are two main streams - continuous diffusion models and discrete diffusion models.
Continuous diffusion models Diffusion-LM (Li et al., 2022) propose a method of mapping tokenized sequences to embedding dimensions guided by a pretrained classifier. DiffuSeq-v1, v2 (Gong et al., 2023a, b) apply partial noising techniques. The core of these diffusion methodologies lies in the addition and restoration of random noise to facilitate generation. However, authors of Cold Diffusion (Bansal et al., 2022) argue that such operations do not necessarily have to be governed by stochastic randomness.
Discrete diffusion models Building on this rationale, D3PM (Austin et al., 2023) propose the discrete restoration-generation approach and DiffusionBERT He et al. (2022) adopt PLMs to DDLM. SEDDLou et al. (2024) propose score entropy inspired by MLM loss. Recent works by Shi et al. (2024) and Sahoo et al. (2024) extend this idea to propose a simplified view of the discrete framework, and obtain better empirical results. Zheng et al. (2024) further close the gap between discrete diffusion models and masked models, and also, by correcting the numerical precision error during the sampling process in SEDD-based models, reveal that the true generation perplexity of DDLMs lags behind that of AR models. These research make an improvement on the open ended generation task. Furthermore, Venkatraman et al. (2024) showcase that SEDD priors can be used to sample from a posterior distribution, namely text infilling, and Nie et al. (2024) demonstrate that DDLMs are scalable, and have advantages over traditional AR models.
8 Conclusions
In this work, we introduce Diffusion-EAGS, integrating PLMs with diffusion models for conditional generation. By leveraging step-wise diffusion models and the property of MLM, we demonstrate that MLM-based PLMs can serve as denoising function with Entropy-based Adaptive Gibbs Sampling. We also propose entropy-based noise scheduling during training for adaptive sampling. Experimental results demonstrate that Diffusion-EAGS outperforms existing baselines, yielding improved text generation quality, enhanced diversity, broad applicability, and more precise token-level control.
Limitations
While Diffusion-EAGS demonstrates significant improvements in conditional generation tasks, there are several limitations. Firstly, as our method is currently focused on text generation tasks, its applicability to text classification tasks, such as Named Entity Recognition and Part-of-Speech Tagging, remains unexplored. Future research could explore extending this method to other NLP tasks. Secondly, future works could investigate the potential in reducing computational costs through more efficient entropy calculation techniques. Lastly, although our current efforts concentrate on developing and validating our framework using encoder-only and encoder-decoder architectures, the potential integration of autoregressive (AR) models remains unexplored. However, since the denoising function—with its entropy-based, sequential ordering—aligns well with autoregressive decoding strategies, we posit that AR models can also serve as effective initializations within our proposed framework.
Ethical Statements
A notable advantage of Diffusion-EAGS is its efficient training process, which optimizes computational resources with high performance. This efficiency not only reduces the environmental impact of training PLMs but also makes advanced NLP technologies more accessible. Additionally, we have focused on output control and interpretability through the use of a discrete diffusion model. Consequently, our methodology contributes to the effective control of generated outputs in the future. Diffusion-EAGS’s strength in tasks involving conditional generation proves beneficial in the field of text generation following social guidelines, where outputs need to strongly depend on the conditions. Specifically, its high diversity performance enables active utilization in dynamically adapting text generations to meet specific social criteria.
Appendix A Measuring Potential Function $\phi(X)$ in MLM
A.1 Experimental Setup
  • Potential Calculation The potential function $\phi(X)$ is computed as described in Equation 1 and  3 , where the log-potentials are obtained for each token using masked token logits.
  • Datasets
    • RocStories: Structured narratives from the RocStories dataset.
A.2 Results of Experiment and Implications for Conditional Generation
The results demonstrate the significant influence of dataset structure on the potential function in MLMs. Specifically, structured datasets like RocStories enforce stronger narrative constraints, leading to higher potentials and greater coherence in sequence generation. This supports the main text’s argument that conditionally constrained spaces enhance the consistency and predictability of MLM outputs.
Appendix B BERT as a MRF
Let $X=(x_{1},\ldots,x_{L})$ be a sequence of random variables $x_{l}$, each taking a value from a vocabulary $V=\{v_{1},\ldots,v_{M}\}$. These variables form a fully-connected graph, indicating mutual dependence.
To define the MRF, we focus on the full-graph clique potential:
$\phi(X)=\exp\left(\sum_{l=1}^{L}\log\phi_{l}(X)\right)$ (3)
where each log-potential $\phi_{l}(X)$ is:
$\log\phi_{l}(X)=$$\begin{cases}1h(x_{l})^{T}f_{\theta}(X\backslash\{x_{l}\}),&\text{if }\text{[MASK]}\notin X_{l-1:l+1}\\ \\ 0,&\text{otherwise}\end{cases}$ (4)
Here, $f_{\theta}(X\backslash\{x_{l}\})$ is a function that maps the sequence $X\backslash\{x_{l}\}$ to a real-valued vector in $\mathbb{R}^{M}$, and $1h(x_{l})$ is a one-hot vector with $x_{l}$ set to 1. from this log potential, we can find probability of sequence X:
$p_{\theta}(X)=\frac{1}{Z(\theta)}\prod_{l=1}^{L}\phi_{l}(X)$
with the normalization constant $Z(\theta)$ defined as:
$Z(\theta)=\sum_{X^{\prime}}\prod_{l=1}^{L}\phi_{l}(X^{\prime})$
For a fixed $X\backslash\{x_{l}\}$, the conditional probability of $x_{l}$ is:
$p(x_{l}|X\backslash\{x_{l}\})=\frac{1}{Z(X\backslash\{x_{l}\})}\exp(1h(x_{l})^{T}f_{\theta}(X\backslash\{x_{l}\}))$
where $Z(X\backslash\{x_{l}\})$ is:
$Z(X\backslash\{x_{l}\})=\sum_{m=1}^{M}\exp(1h(m)^{T}f_{\theta}(X\backslash\{x_{l}\}))$
BERT uses pseudo log-likelihood learning to handle the intractability of $Z(\theta)$:
$\text{PLL}(\theta;D)=\frac{1}{|D|}\sum_{X\in D}\sum_{l=1}^{|X|}\log p(x_{l}|X\backslash\{x_{l}\})$
where $D$ is the training dataset.
Appendix C Dataset Explanations
Quasar-T QQP ParaDetox Deontology RocStories
input output input output input output input output input output
Max 63 244 104 98 35 35 24 31 76 57
Table 8: Dataset Statistics
Appendix D Detailed analysis of Results
D.1 Ours vs AR model
Notably, our model effectively reflects the pattern and style of the dataset, attributed to the capability of encoder-based PLMs. Specifically, as reported in Table 2-ParaDetox, our model excels at capturing semantic information with the highest MAUVE score of 0.733, despite ParaDetox being a challenging colloquial dataset including slang, numerous abbreviations, and various perturbations. In addition, our model demonstrate robustness to such perturbations with 69.5 PPL, while GPT-2 shows a lower performance of 389.1 PPL. Diffusion-EAGS demonstrates high level of text quality surpassing that of GPT-2 in Table 3 in text quality.
As for diversity, our model consistently outperforms GPT models in VS(ngram) and VS(emb) in Table 23 and 4. These results underscore our model’s efficacy in generating diverse responses. Specifically, in the context of story generation as shown in Table 4, our model not only demonstrates effectiveness in enhancing both the quality and diversity of the text but also maintains coherence with the contextual storyline. This is further supported by the high LLM-c scores, which confirm the plausibility of the narratives produced by our model.
D.2 Ours vs CDLMs
Notably, CDLMs demonstrate a noticeable deficiency in diversity. In contrast, our model excels at producing significantly more diverse sequences. This enhanced performance can be attributed to the integration of PLM into our framework, which evidently enriches the diversity of the generated outputs. In addition, our model generally achieves higher text quality scores than CDLMs across all experiments. In case, our model shows a lower PPL than DINOISER, but DINOISER’s MAUVE score approaches zero, indicating that their generated responses are not based on the dataset. Diffusion-EAGS not only surpasses Diffuseq and DINOISER in PPL and SOME but also achieves a MAUVE score improvement of over 0.4, thus indicating that the generated responses accurately reflect the dataset while preserving grammatical integrity. Furthermore, our models require only a few steps, resulting in a 40% reduction of costs with higher quality and diversity.
D.3 Ours vs DDLMs
Contrary to CDLMs, DiffusionBERT exhibits limitations in text quality, as shown in Table 234. On the other hand, our model achieves significantly higher scores across all quality metrics. This observation suggests that the naive application of BERT (Devlin et al., 2019) within the diffusion process fails to fully harness the capabilities of PLMs. SEDD, an open-ended generation model shows low performance under text generation in highly constrained tasks such as ParaDetox. Even in open-ended generation settings of Table 4, Diffusion-EAGS shows higher text quality and semantic scores. Additionally, our model does not require more than 5 diffusion steps, as it can adaptively recover the absorbing state through an entropy-based generation approach. Above results further exemplify the high performance and the time efficiency of our model.
Appendix E Experimental Outputs
E.1 LLM Evaluation
Table 9: ParaDetox LLM Evalution (LLM-t)

### v3

@@ F.2 Dataset Explanations
F.3 Experimental Details
@@
We employ roberta-base as MLM with learning rate 5e-4. The maximum lengths for QG, QQP, and Paradetox is set to 64, while for Deontology and DialogSum set to 48 and 292, respectively, based on data statistics. We test 20 conditions with 5 outputs in total 100, which is not used for training. The number of steps is configured to 5. We then perform a naive categorical sampling with a sample size of 20 and select final 5 samples based on PPL. We use 1 A100 GPU with the batch size as 256. Experimental details of LLMs are in Appendix J, machine translation in Appendix H.
Quality metrics
To measure the quality of the generated texts, we use Perplexity based-on GPT-2 Large and GPT-2 XL, SOME (Yoshimura et al., 2020), the grammar metric based on corpus, LLM-c (Lin and Chen, 2023) to measure the plausibility of the narratives, LLM-t (Koh et al., 2024a) to measure toxicity, and MAUVE Pillutla et al. (2021), measuring a reflectiveness of training dataset characteristics of generate outputs. MAUVE score of 1 indicates that the output perfectly matches the training dataset as a neural database. For Mean Opinion Score (MOS), we get 5 outputs from each condition. Subsequently, hired four integrated ph.d course work annotators in the university NLP research lab evaluate the generated text based on two criteria: (1) semantic reflectiveness of the condition, indicating how accurately the condition is represented in the text, and (2) sentence completeness, assessing overall grammatical and semantic coherence. Each criterion was rated on a scale from 0 to 1. Subsequently, these scores are normalized and averaged to obtain a final score ranging from 0 to 1. In our evaluation, Fleiss’ kappa Fleiss (1971) is exceeded 0.7 as assessing sentence quality is both intuitive and relatively non-controversial among the annotators.
Diversity Metrics
Traditional diversity metrics Self-BLEU Zhu et al. (2018) and distinct-n Li et al. (2015) are employed to evaluate the generated texts. We also adopt Vendi Score (VS)-SimCSE Friedman and Dieng (2023), an interpretable diversity metric, which quantifies the effective number of unique samples in a given set. Both the n-gram and embedding variations are utilized, where embedding VS is semantic diversity. For the diversity MOS evaluation, we adopt the same methodology used for the quality MOS but apply two distinct criteria: (1) the condition’s semantic reflectiveness, and (2) sentence diversity, capturing both semantic and structural variety beyond mere word deletion or rearrangement. Ideal score of diversity MOS is 5 which means different five sequences for one condition, and lowest score is 1 which means all identical sequences.
@@
Appendix G Detailed analysis of Results
G.1 Fine-Grained Comparison
As shown in Table 289, our model consistently exhibits exceptional performance in terms of text quality while simultaneously maintaining diversity when compared to baseline models. The standard deviation of PPL in Paradetox Experiment is 61 for our model. All other PPL’s standard deviation are similar to that of Paradetox.
In Table 8 Paradetox, our model demonstrates superior performance across all evaluated metrics. Such phenomenon represents that our model based on MLM shows robustness on diverse perturbations of daily dialogues. When PPL exceeds 600, the model is considered to have failed in generating natural sequences, and is thus represented in gray color. Specifically, the text quality produced by the CMLM, which is standard BERT-generation, and SEDD, which is powerful model in open-ended generation, is found to be low.
Consequently, these models were excluded from subsequent experiments. In Deontology, our model exceeds the baseline models’ PPL and MAUVE scores, whereas SOME score represent the sufficient quality of text with the highest diversity score. As illustrated in Table 9, Diffusion-EAGS generates the responses with the highest PPL score for QG, and highest MAUVE and PPL score for QQP.
While we adhere to the standard metrics commonly used in diffusion research and integrate as many additional metrics as possible, we also comprehensively explore our model’s capabilities across multiple dimensions. As the outputs of earlier generation tasks are too broad to be effectively evaluated using reference-based metrics, we provide generated examples in Appendix I and measure the preference of these outputs using a LLM-based metric in Appendix G.2. Additionally, to accommodate a scenario where reference-based evaluation is applicable, we have included a more extensive summarization task in Appendix G.2 and translation task in Appendix G.3. These results confirm that our method consistently produces outputs that adhere to the specified conditions.
Diffusion-EAGS demonstrates the highest MAUVE score of 0.733 in Table 8-ParaDetox, and high level of text quality surpassing that of GPT-2 in Table 9 in text quality. ParaDetox is colloquial dataset including slang, numerous abbreviations, and various perturbations, so our model demonstrate robustness to such perturbations with 69.5 PPL. As for diversity, our model consistently outperforms GPT models in VS(ngram) and VS(emb) in Table 28, and 9.
@@ G.1 Fine-Grained Comparison
G.2 Quality Recheck – LLM score & Dialogue Summarization
Model ROUGE-1 ROUGE-2 MAUVE Ngram Emb Self-BLEU Distinct-1 Distinct-2
Ours 0.409 0.174 0.536 4.114 2.591 0.252 0.253 0.632
SEDD 0.179 0.032 0.999 4.216 2.576 0.211 0.200 0.609
DINOISER 0.209 0.031 0.337 1.247 1.227 0.926 0.256 0.633
Table 10: DialogueSum Experiment
Model PPL MAUVE vs(ngram) VS(emb) sef-bleu distinct-1 distinct-2
GENIE 134.1 0.296 2.527 1.800 0.702 0.454 0.825
Table 11: Quantitative results for the GENIE model.
Paradetox w/ LLM-t on application models
LLM-t
GPT-2 0.02
GPT-3.5 0.074
GPT-4 0.18
DiffuSeq 0.03
Diffusion-Bert 0.09
DINOISER 0.1
SEDD-small NA
Diffusion-EAGS 0.01
Table 12: ParaDetox Dataset Generation – LLM-t is the LLM-evaluation for measuring toxicity.
Since our research primarily aims to enhance the model’s inherent capabilities, we set up baselines that revolve around (or are closely related to) noise scheduling. Nevertheless, some studies employ a hybrid framework integrating LLMs and diffusion models Lin et al. (2023); Xiang et al. (2024); hence, we conduct additional experiments to investigate this scenario. In addition, to evaluate the quality of the paradetox output and ours diffusion-EAGS still outperforms GENIE (Lin et al., 2023) in Table 11. We also use the LLM-t score Koh et al. (2024b) to measure whether models successfully detoxify the source condition, showing the quality of generated outputs from ours as shown in Table 12.
Models Prefer Baseline Prefer Ours Tie
diffuseq vs. ours 20% 65% 15%
diffusionBERT vs. ours 20% 65% 15%
dinoiser vs. ours 0% 90% 10%
GPT-2 vs. ours 25% 65% 10%
Table 13: Evaluation results comparing our model with various baselines.
QG - LLM preference
For Question Generation (QG), we employ the widely adopted GPT-as-a-Judge framework Zheng et al. (2023) to evaluate the quality of generations produced by our model and the baselines on the QG dataset. We adopt a pairwise evaluation setting, following the system and input prompts specified in Zheng et al. (2023) for the pairwise comparison. The factors specified to be evaluated are 1) coherency 2) grammatical correctness 3) semantic soundness 4) diversity and 5) being a more reasonable question to the input (condition) text. We employ the gpt-4 model. The result is as follows:
Note that since within the prompt, the baseline model’s generations are specified prior to our model’s generation, there is a significant position bias working against our favor, as noted in Zheng et al. (2023). The results above indicate that despite such bias, our model’s generations are much more favored over the baselines’ generations.
Dialoguesum Experiment
Our model outperforms existing baselines in ROUGE, a reference-based metric as shown in Table 10. These findings indicate that, according to the automatic scores, our model sufficiently captures the source condition.
Human Evaluation
Below, we report the Mean Opinion Score (MOS) averages and standard deviations (std) in the following order: DiffusionBERT, LD4LG, GPT-2, Dinoiser, and our method. First, the average scores of semantic reflection are 0.98, 0.90, 0.94, 0.98, and 0.97, respectively, with standard deviations of 0.14, 0.30, 0.24, 0.14, and 0.16. Second, the average scores of sentence completeness are 0.78, 0.92, 0.72, 0.84, and 0.90, respectively, with standard deviations of 0.18, 0.14, 0.28, 0.15, and 0.15. Third, average scores of diversity are 2, 1, 2.65, 1, and 4.6, respectively, with standard deviations of 1.3, 0, 1.45, 0, and 0.7. GPT-3.5-turbo’s std is almost 0.
@@ G.2 Quality Recheck – LLM score & Dialogue Summarization
G.3 Machine Translation : Bilinguality & Low Resource Settings
Labeled datasets used in conditional generation tasks are typically limited in size and sometimes multilingual. To further assess our model’s performance in conditional generation, particularly in terms of language extension and resource scarcity, we conduct additional experiments on a translation task. We conduct additional experiments on CMLMs such as Mask-and-Predict and Easy-First, diffusion models such as Diffuseq-v2 (Gong et al., 2023b) and SEDD, traditional translation models such as mBART-50 (Tang et al., 2020) and NLLB. For evaluation metrics, we utilize sacreBLEU (Post, 2018) and neural-net scores such as COMET (Rei et al., 2020) and XCOMET (Guerreiro et al., 2023). More details are provided in Appendix L.
Table 14 shows that predicting the target sequence without leveraging a multilingual model proves to be challenging. All diffusion baseline models struggle to produce correct outputs. Similar challenges arise in NAR transformer baselines. Despite constructing the vocabulary using the pretrained mBART-50 model (DisCo-m), the underlying issues remain. On the other hand, our proposed model demonstrates promising results.
G.4 Diversity Analysis
Limitation of Diversity on Traditional DDLMs
GPT-2 Given Source : holy shit , they blew up a real artifact this time ? from GPT2s output ended by end token with default temperature sampling from huggingface : - Oh my god,they blew a really important artifact in this year?????..safe: Oh - Oh my god, they destroyed a really important artifact in this year?...safe: Oh God,they - they blow up something thistime?.??.safe: Oh my god, they destroyed a really valuable artifact - Oh my god, they destroyed a really important artifact in this year??.?!??.?!? - They blew an artifact that time? Theyre still in the artifact? This time, theyre in trouble. This
Tranditional Diffusion Models traditional diffusion models output from Dinoiser, LD4LG : - they blew up a real artifact this time? - they blew up a artifact this time? - they blew up a real artifact? - they blew up a real artifact this time? - they blew up a real artifact this time?
Ours : Diffusion-EAGS from ours : - aww, it is really a real artifact this time ? - it seems like they destroyed an artifact in this time as well - they have blown up a large artifact - they have blown up it in a museum, this time - they also destroyed artifacts at the same time
We summarize the generation trends of the models presented in Table above. We observe that when a fine-tuned GPT-2 is tasked with strongly constrained conditional generation, it struggles to properly terminate sentences with an <eos> token. In particular, it shows limitations when handling semantic leaps or clearly delineated structural constraints, leading to suboptimal conditional generations.
Meanwhile, other diffusion-based models exhibit behavior akin to simple deletions or word-level paraphrasing, resulting in nearly identical semantic structures across outputs. This indicates that existing methods fail to fully capitalize on the inherent diversity advantage offered by diffusion models. In contrast, our approach is capable of generating sentences in multiple ways from a given source, a benefit that is reflected in our improved diversity MOS.
Diversity Saturation on LLMs
@@ G.4 Diversity Analysis
Inspired by the observation that Diffusion-EAGS consistently excel in terms of diversity across all results, we delve further into the diversity capabilities of our model. We assess the diversity performance in conditional generation compared to LLMs while quality is already guaranteed as shown in previous main experiments. We measure the VS for 5 to 100 generations under a single condition. Such experiment demonstrates the extent to which the model’s output diversity saturates, enabling a comparison of asymptotic diversity performance. The experiment is conducted on the ‘deontology’ dataset which allows high output diversity in its settings. Details of using LLMs are provided in Appendix J.
@@
G.5 Keyword Generation Results
As shown in Table 4, our model successfully generate the coherent sequences based on givel context and keywords.
G.6 Different PLM : BERT & T5
While our primary approach integrates BERT into the diffusion framework via a theoretical cMRF interpretation, we also experiment with other PLMs such as BERT and T5, because the main experiment involved BART (LD4LG) and GPT-2. Specifically, RoBERTa and BERT exhibit similar trends, whereas T5 shows behavior comparable to a fine-tuned GPT-2 in Table 15. We conjecture that T5 is already trained with an autoregressive strategy in its decoder whose generation process is largely influenced by its initial decoder tokens from an entropy perspective Wang and Zhou (2024), resulting in relatively low diversity. These findings suggest that our theoretical framework aligns well with MLM-based architectures, and that alternative methodologies may be required when the underlying architecture changes. Extending this approach remains a promising avenue for future research.
Model PPL MAUVE VS(ngram) vs(emb) self-bleu distinct-1 distinct-2
RoBERTa 109.3 0.811 4.417 3.311 0.256 0.407 0.810
BERT 69.5 0.773 4.755 3.659 0.126 0.475 0.834
T5 408.1 0.378 2.256 1.666 0.750 0.415 0.773
Table 15: Performance comparison of T5, BERT, and RoBERTa.
Appendix H Experimental Outputs
H.1 LLM Evaluation
The LLM evaluation prompt for ParaDetox is provided in Table 16, and the LLM evaluation prompt for RocStories is given in Table 17.